OpenStudy (anonymous):

what is the equation of the line passing through (-6,1) and is perpendicular to the line y=-3x+1 in slope intercept form . I keep getting stuck halfway through so work would be helpful

OpenStudy (anonymous):

the slope would change to 1/3

OpenStudy (anonymous):

Answer choices: y=3x+4 y=1/3x+3 y=-3x-6 y=-1/3x-5 would it be the second choice?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

any time!